# REPS: Recycling Entropies for Packet Spraying to Adaptively Explore Paths and Mitigate Failures

### v1

@@ Abstract.
Most existing datacenter transport protocols rely on in-order packet delivery, a design choice rooted in legacy systems and simplicity. However, advancements in technology, such as RDMA, have made it feasible to relax this requirement, allowing for more effective use of modern datacenter topologies like FatTree and Dragonfly. The rise of AI/ML workloads underscores the necessity for enhanced link utilization, a challenge for single-path load balancers due to issues like ECMP collisions.
In this paper, we introduce REPS, a novel per-packet traffic load-balancing algorithm that integrates seamlessly with existing congestion control mechanisms. REPS reroutes packets around congested hotspots and unreliable or failing links with remarkable simplicity and minimal state requirements.
Our evaluation demonstrates that REPS significantly outperforms traditional packet spraying and other state-of-the-art solutions in contemporary datacenter networks, offering substantial improvements in performance and link utilization.
@@ 1. Introduction
For this reasons, we propose REPS, a simple load balancing scheme based on adaptive packet spraying. We design REPS in order to work with any congestion control scheme and transport layer as long as it is able to handle out-of-order packets which is what next generation protocols are pushing for given the inherit performance advantage.
REPS does not require any specific hardware support from switches beside ECMP which is commonly present in modern switches. If available, REPS can take advantage of packet trimming to further improve its accuracy but it is also capable of running without that.
@@ 1.2. Contribution
- REPS:
A powerful and lightweight adaptive load balancing scheme called Recycled Entropy Packet Spraying (REPS, Sec.3). REPS achieves near-optimal performance at high bandwidth utilization and efficiently routes around path failures, congestion hotspots, and network asymmetries. Unlike more complex in-network load balancing solutions(De Sensi et al., 2019; Rocher-González et al., 2022), REPS does not require any switch support beyond ECMP. REPS provides a significant improvement over traditional packet spraying, both in perfectly symmetric networks and in the presence of asymmetries, whether due to differing topologies or potential failures. We demonstrate that REPS performs well regardless of the congestion control scheme used.
@@ 2.1. Congestion Signals
In this section we are going over several congestion signals that REPS can take advantage of if they are implemented in the network. At minimum we require ECN to be supported by the switches.
@@
REPS utilizes ECN as its primary congestion signal to determine if a path is congested. Notably, ECN marking is performed at the egress to ensure that the most current information is captured. If packet trimming is supported, REPS leverages this feature to enhance its ability to distinguish between packets lost due to congestion and those lost due to network failures. In situations where packet trimming is not supported, REPS employs a straightforward timeout mechanism to detect packet loss within the network.
3. Recycled Entropy Packet Spray (REPS)
REPS is a simple load balancer scheme that does not need any switch support besides ECMP (Hopps, 2009). With ECMP, each switch computes a hash function over some packet header fields to determine which path to forward the packet on. One of these fields is the entropy (e.g., the source port or the IPv6 Flow Label), which REPS changes on a per-packet basis.
The key idea behind REPS is very simple: we want to explore new possible paths when congestion is detected while re-using paths where there is little to no congestion. This has the immediate advantage of being both simple and lightway since we can potentially store only the last good entropy that we have seen and which is to be reused. However, the slightly more advanced version of it uses a circular buffer of a fixed size where we save all the cached entropies.
For the first of packets, REPS explores all possible entropies. This is necessary since initially we do not have any information about the status of the network. The receiver copies the entropies from the received packet to the ACK, thus forwarding them back to the sender. Whenever an ACK arrives, if the ACK is not ECN-marked, the entropy carried by the ACK is cached. Intuitively, because the corresponding data packet did not experience congestion, we can send other packets on the same path, and REPS will thus reuse it for subsequent packets. Otherwise, if the ACK was ECN-marked, REPS selects the next entropy. This can be further extended by defining additional policies (e.g., excluding bad entropies after some time or storing entropies in a priority queue based on their RTTs and/or ECN marking), but we do not explore this further in this initial version of this work. The circular buffer design guarantees that even when receiving back-to-back ACKs with good entropies, these can be correctly stored and used. On the sending side, if we have multiple elements in the buffer, we will always first select the oldest entropy to make sure to not get stuck on old information.
We evaluate REPS’ advantages versus traditional packet spraying and ECMP in Sec. 4.2. Algorithm 1 describes the REPS pseudocode, and Fig. 1(a) shows an example of REPS’ behavior.
@@ 3.0.1. Interaction between CC and load balancing
We design REPS with the idea of it being capable of working with any transport layer that does not require in-order delivery of packets. Having said that, co-designing the congestion control protocol with the load balancer algorithm can bring several benefits. For these reasons, we optimize REPS to work with SMaRTT (Bonato et al., 2024). REPS effectively mitigates the effects of fabric congestion, such as those caused by ECMP hashing collisions, without necessitating a reduction in the congestion window size. This capability is particularly important in permutation workloads on non-oversubscribed networks, where proper traffic balancing can theoretically prevent congestion. To enable REPS to respond to such congestion, SMaRTT is designed to ignore individual ECN-marked packets unless multiple marked packets are received. Specifically, SMaRTT maintains an exponentially weighted moving average of ECN-marked packets, updating the variable using the formula . The congestion window remains unchanged as long as , meaning that the system does not react to congestion unless 25 percent or more of the acknowledgments indicate congestion—a threshold high enough to exclude transient load imbalances. This mechanism, referred to as Wait to Decrease (WTD), is illustrated in Fig. 1, which demonstrates how it significantly improves the flow completion time (FCT) for permutation workloads running on a 1,024-node non-oversubscribed fat tree network.
@@ 4. Evaluation
We pair REPS together with SMaRTT and compare it against packet spraying and ECMP. We use SMaRTT for all of them in order to have the same baseline regarding congestion control. We enable trimming for all the algorithms.
@@ 4.1. Non-Oversubscribed Networks
In this section we evaluate the performance of REPS in a simple setting where there is no ovserubscription and no failures, meaning the network is perfectly symmetrical. Intuitively, this is the best situation for packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we can see in Fig.2(a), this is not the case as REPS still offers a 10% advantage over oblivious packet spraying. This is because of ECMP collisions that still happen with spraying. While over long period of time, each link will be evenly used, there will still be short term collisions happening that also increase the RTT of packets as seen in Fig.2(b).
@@ 4.2. Oversubscribed Networks
In Fig.3(a), we evaluate the performance of SMaRTT with various load balancing algorithms: REPS, ECMP(Hopps, 2009), and oblivious packet spraying (Dixit et al., 2013). We simulate a 4 MiB permutation workload on a 8:1 oversubscribed fat tree network. The results show that REPS achieves again approximately a 10% improvement in completion time over oblivious packet spraying due to its superior fairness. In contrast, per-flow ECMP results in job completion times that are 1.5 times higher than those achieved with packet spraying and REPS, primarily because of increased flow collisions (which are not depicted in the plot). However, the individual RTTs for per-flow ECMP are relatively low since SMaRTT reduces the congestion window size upon detecting congestion caused by hash collisions. Additionally, an analysis of the packet RTT distribution in Fig. 3(b) reveals that REPS maintains RTTs that are closer to the target with lower variability, indicating improved network utilization.
@@ 4.3. Failures
We now evaluate what happens in a network where some links are failing, in particular we fail a fixed number of links between the aggregation switches and core switches of the fat tree. REPS now gains a 50% advantage (Fig. 4(b)) over spraying because it is able to avoid more efficiently the failing links, while spraying would randomly select failing paths over time and hence increase the completion time. For ECMP, certain flows don’t properly terminate as they are not able to adaptively avoid failing links.
@@ 5. Conclusion
We design REPS, a simple and resource-efficient load balancer that can be paired with different congestion control algorithms. REPS significantly improves the FCT, runtime and number of dropped packets as shown in our evaluation. REPS can drastically improve the load balancing performance compared to traditional per-flow ECMP or oblivious packet spraying at minimal complexity cost. We also show that REPS can effectively handle faults and network asymmetries. In a future version of this work we will expand REPS to be used alongside several congestion control algorithms and with different failure modes.

### v2

# ARCANE: Adaptive Routing with Caching and Network Exploration

@@ Abstract.
Most datacenter transport protocols traditionally depend on in-order packet delivery, a legacy design choice that prioritizes simplicity. However, technological advancements, such as RDMA, now enable the relaxation of this requirement, allowing for more efficient utilization of modern datacenter topologies like FatTree and Dragonfly. With the growing prevalence of AI/ML workloads, the demand for improved link utilization has intensified, creating challenges for single-path load balancers due to problems like ECMP collisions.
In this paper, we present ARCANE, a novel, adaptive per-packet traffic load-balancing algorithm designed to work seamlessly with existing congestion control mechanisms. ARCANE dynamically routes packets to bypass congested areas and network failures, all while maintaining a lightweight footprint with minimal state requirements.
Our evaluation shows that ARCANE delivers significant performance gains over traditional load-balancing methods, including packet spraying and other advanced solutions, substantially enhancing both performance and link utilization in modern datacenter networks.
@@ 1. Introduction
For this reasons, we propose ARCANE, a simple load balancing scheme based on adaptive packet spraying. We design ARCANE in order to work with any congestion control scheme and transport layer as long as it is able to handle out-of-order packets which is what next generation protocols are pushing for given the inherit performance advantage.
ARCANE does not require any specific hardware support from switches beside ECMP which is commonly present in modern switches. If available, ARCANE can take advantage of packet trimming to further improve its accuracy but it is also capable of running without that.
@@ 1.2. Contribution
- ARCANE:
A powerful and lightweight adaptive load balancing scheme called Adaptive Routing with Caching and Network Exploration (ARCANE, Sec.3). ARCANE achieves near-optimal performance at high bandwidth utilization and efficiently routes around path failures, congestion hotspots, and network asymmetries. Unlike more complex in-network load balancing solutions(De Sensi et al., 2019; Rocher-González et al., 2022), ARCANE does not require any switch support beyond ECMP. ARCANE provides a significant improvement over traditional packet spraying, both in perfectly symmetric networks and in the presence of asymmetries, whether due to differing topologies or potential failures. We demonstrate that ARCANE performs well regardless of the congestion control scheme used.
@@ 2.1. Congestion Signals
In this section we are going over several congestion signals that ARCANE can take advantage of if they are implemented in the network. At minimum we require ECN to be supported by the switches.
@@
ARCANE utilizes ECN as its primary congestion signal to determine if a path is congested. Notably, ECN marking is performed at the egress to ensure that the most current information is captured. If packet trimming is supported, ARCANE leverages this feature to enhance its ability to distinguish between packets lost due to congestion and those lost due to network failures. In situations where packet trimming is not supported, ARCANE employs a straightforward timeout mechanism to detect packet loss within the network.
3. Adaptive Routing with Caching and Network Exploration (ARCANE)
ARCANE is a simple load balancer scheme that does not need any switch support besides ECMP (Hopps, 2009). With ECMP, each switch computes a hash function over some packet header fields to determine which path to forward the packet on. One of these fields is the entropy (e.g., the source port or the IPv6 Flow Label), which ARCANE changes on a per-packet basis.
The key idea behind ARCANE is very simple: we want to explore new possible paths when congestion is detected while re-using paths where there is little to no congestion. This has the immediate advantage of being both simple and lightway since we can potentially store only the last good entropy that we have seen and which is to be reused. However, the slightly more advanced version of it uses a circular buffer of a fixed size where we save all the cached entropies.
For the first of packets, ARCANE explores all possible entropies. This is necessary since initially we do not have any information about the status of the network. The receiver copies the entropies from the received packet to the ACK, thus forwarding them back to the sender. Whenever an ACK arrives, if the ACK is not ECN-marked, the entropy carried by the ACK is cached. Intuitively, because the corresponding data packet did not experience congestion, we can send other packets on the same path, and ARCANE will thus reuse it for subsequent packets. Otherwise, if the ACK was ECN-marked, ARCANE selects the next entropy. This can be further extended by defining additional policies (e.g., excluding bad entropies after some time or storing entropies in a priority queue based on their RTTs and/or ECN marking), but we do not explore this further in this initial version of this work. The circular buffer design guarantees that even when receiving back-to-back ACKs with good entropies, these can be correctly stored and used. On the sending side, if we have multiple elements in the buffer, we will always first select the oldest entropy to make sure to not get stuck on old information.
We evaluate ARCANE’ advantages versus traditional packet spraying and ECMP in Sec. 4.2. Algorithm 1 describes the ARCANE pseudocode, and Fig. 1(a) shows an example of ARCANE’ behavior.
@@ 3.0.1. Interaction between CC and load balancing
We design ARCANE with the idea of it being capable of working with any transport layer that does not require in-order delivery of packets. Having said that, co-designing the congestion control protocol with the load balancer algorithm can bring several benefits. For these reasons, we optimize ARCANE to work with FASTFLOW (bonato2024FASTFLOWARCANEsenderbasedmarkedrapidlyadapting). ARCANE effectively mitigates the effects of fabric congestion, such as those caused by ECMP hashing collisions, without necessitating a reduction in the congestion window size. This capability is particularly important in permutation workloads on non-oversubscribed networks, where proper traffic balancing can theoretically prevent congestion. To enable ARCANE to respond to such congestion, FASTFLOW is designed to ignore individual ECN-marked packets unless multiple marked packets are received. Specifically, FASTFLOW maintains an exponentially weighted moving average of ECN-marked packets, updating the variable using the formula . The congestion window remains unchanged as long as , meaning that the system does not react to congestion unless 25 percent or more of the acknowledgments indicate congestion—a threshold high enough to exclude transient load imbalances. This mechanism, referred to as Wait to Decrease (WTD), is illustrated in Fig. 1, which demonstrates how it significantly improves the flow completion time (FCT) for permutation workloads running on a 1,024-node non-oversubscribed fat tree network.
@@ 4. Evaluation
We pair ARCANE together with FASTFLOW and compare it against packet spraying and ECMP. We use FASTFLOW for all of them in order to have the same baseline regarding congestion control. We enable trimming for all the algorithms.
@@ 4.1. Non-Oversubscribed Networks
In this section we evaluate the performance of ARCANE in a simple setting where there is no ovserubscription and no failures, meaning the network is perfectly symmetrical. Intuitively, this is the best situation for packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we can see in Fig.2(a), this is not the case as ARCANE still offers a 10% advantage over oblivious packet spraying. This is because of ECMP collisions that still happen with spraying. While over long period of time, each link will be evenly used, there will still be short term collisions happening that also increase the RTT of packets as seen in Fig.2(b).
@@ 4.2. Oversubscribed Networks
In Fig.3(a), we evaluate the performance of FASTFLOW with various load balancing algorithms: ARCANE, ECMP(Hopps, 2009), and oblivious packet spraying (Dixit et al., 2013). We simulate a 4 MiB permutation workload on a 8:1 oversubscribed fat tree network. The results show that ARCANE achieves again approximately a 10% improvement in completion time over oblivious packet spraying due to its superior fairness. In contrast, per-flow ECMP results in job completion times that are 1.5 times higher than those achieved with packet spraying and ARCANE, primarily because of increased flow collisions (which are not depicted in the plot). However, the individual RTTs for per-flow ECMP are relatively low since FASTFLOW reduces the congestion window size upon detecting congestion caused by hash collisions. Additionally, an analysis of the packet RTT distribution in Fig. 3(b) reveals that ARCANE maintains RTTs that are closer to the target with lower variability, indicating improved network utilization.
@@ 4.3. Failures
We now evaluate what happens in a network where some links are failing, in particular we fail a fixed number of links between the aggregation switches and core switches of the fat tree. ARCANE now gains a 50% advantage (Fig. 4(b)) over spraying because it is able to avoid more efficiently the failing links, while spraying would randomly select failing paths over time and hence increase the completion time. For ECMP, certain flows don’t properly terminate as they are not able to adaptively avoid failing links.
@@ 5. Conclusion
We design ARCANE, a simple and resource-efficient load balancer that can be paired with different congestion control algorithms. ARCANE significantly improves the FCT, runtime and number of dropped packets as shown in our evaluation. ARCANE can drastically improve the load balancing performance compared to traditional per-flow ECMP or oblivious packet spraying at minimal complexity cost. We also show that ARCANE can effectively handle faults and network asymmetries. In a future version of this work we will expand ARCANE to be used alongside several congestion control algorithms and with different failure modes.